### v2

# A note on the Erdős Matching Conjecture

Abstract.
The Erdős Matching Conjecture states that the maximum size $f(n,k,s)$ of a family ${\mathcal{F}}\subseteq\binom{[n]}{k}$ that does not contain $s$ pairwise disjoint sets is $\max\{|{\mathcal{A}}_{k,s}|,|{\mathcal{B}}_{n,k,s}|\}$, where ${\mathcal{A}}_{k,s}=\binom{[sk-1]}{k}$ and ${\mathcal{B}}_{n,k,s}=\{B\in\binom{[n]}{k}:B\cap[s-1]\neq\emptyset\}$. The case $s=2$ is simply the Erdős-Ko-Rado theorem on intersecting families and is well understood. The case $n=sk$ was settled by Kleitman and the uniqueness of the extremal construction was obtained by Frankl. Most results in this area show that if $k,s$ are fixed and $n$ is large enough, then the conjecture holds true. Exceptions are due to Frankl who proved the conjecture and considered variants for $n\in[sk,sk+c_{s,k}]$ if $s$ is large enough compared to $k$. A recent manuscript by Guo and Lu considers non-trivial families with matching number at most $s$ in a similar range of parameters.
In this short note, we are concerned with the case $s\geq 3$ fixed, $k$ tending to infinity and $n\in\{sk,sk+1\}$. For $n=sk$, we show the stability of the unique extremal construction of size $\binom{sk-1}{k}=\frac{s-1}{s}\binom{sk}{k}$ with respect to minimal degree. As a consequence we derive $\lim\limits_{k\rightarrow\infty}\frac{f(sk+1,k,s)}{\binom{sk+1}{k}}<\frac{s-1}{s}-\varepsilon_{s}$ for some positive constant $\varepsilon_{s}$ which depends only on $s$.
Research partially supported by Simons Foundation Collaboration Grant for Mathematicians #709641
Research supported by NKFIH under grant FK132060
1. Introduction
We use standard notation. Let $[n]$ denote the set of the first $n$ positive integers. For a set $X$, $\binom{X}{k}$ stands for the family of all $k$-element subsets of $X$ and a $k$-graph (or $k$-uniform hypergraph) on $X$ is a subset of $\binom{X}{k}$. The members of $X$ are called vertices and the $k$-sets are called $k$-edges or hyperedges or just edges. A $k$-graph is intersecting if every pair of edges has a nonempty intersection. A $k$-graph on $X$ for which there is an $x\in X$ that is in every edge is called a star.
One of the major unsolved problems in extremal set theory is due to Erdős: The matching number $\nu({\mathcal{F}})$ of a family ${\mathcal{F}}$ of sets is the size of the largest set of pairwise disjoint sets that ${\mathcal{F}}$ contains. Let $f(n,k,s)$ denote the maximum size of a $k$-graph on $n$ vertices with matching number strictly smaller than $s$. Because $\nu({\mathcal{F}})=1$ if and only if ${\mathcal{F}}$ is intersecting, the well-known Erdős-Ko-Rado theorem [5] gives
$f(n,k,2)=\begin{cases}\binom{n-1}{k-1},&\mbox{ if $n\geq 2k$;}\\ \binom{n}{k},&\mbox{ if $n\leq 2k-1$.}\end{cases}$
The special case of $f(sk,k,s)$ was proved by Kleitman.
Theorem 1 (Kleitman [22]).
If $k\geq 2$ and $s\geq 2$ are positive integers, then $f(sk,k,s)=\binom{sk}{k}-\binom{sk-1}{k-1}=\frac{s-1}{s}\binom{sk}{k}$ with equality if and only if there is an element $x\in[n]$ such that every $k$-edge fails to contain $x$.
There are two natural families of $k$-sets with matching number $s-1$:
${\mathcal{A}}_{k,s}=\binom{[sk-1]}{k}\hskip 28.45274pt{\mathcal{B}}_{n,k,s}=\left\{B\in\binom{[n]}{k}:B\cap[s-1]\neq\emptyset\right\}.$
Erdős conjectured that for any values of $n,k,s$, one of these families achieves the maximum value.
Conjecture 2 (Erdős Matching Conjecture (EMC) [4]).
For all $n,k,s$ with $n\geq sk$, we have $f(n,k,s)=\max\{|{\mathcal{A}}_{k,s}|,|{\mathcal{B}}_{n,k,s}|\}$.
It can be computed that $|{\mathcal{A}}_{k,s}|<|{\mathcal{B}}_{n,k,s}|$ for all $n\geq s(k+1)$. Most results concerning Conjecture 2 determine that $f(n,k,s)=|{\mathcal{B}}_{n,k,s}|$ if $n$ is large enough compared to $k$ and $s$ [15, 20], with the current best bounds due to Frankl [8] for $n\geq(2s+1)k-s$ and Frankl and Kupavskii [14] for $n\geq\frac{5}{3}sk-\frac{2}{3}s$ if $s$ is large enough. The case $k=3$ is also completely settled by Frankl [9]. Minimum degree versions of the problem are also studied [1, 18, 21] and general inequalities on $f(n,k,s)$ are known [14, 19] for $n=cks$ with $c>1$ and $k$ and/or $s$ is large enough.
Not too many results are known in the region when $n$ is very close to $sk$: Frankl [10] proved that $f(n,k,s)=|{\mathcal{A}}_{k,s}|$, if $sk\leq n\leq s(k+\varepsilon)$ where $\varepsilon$ depends only on $k$, but $s$ has to be large enough with respect to $k$ ($s\geq k^{2}+k+1$ suffices). This result was extended by Kolupaev and Kupavskii [23]. Frankl [12] and recently Guo and Lu [17] considered non-trivial families (i.e. those with no isolated vertices) in a similar range of parameters.
Our aim is to obtain bounds on $f(n,k,s)$ when $n=sk+1$, but with $s$ fixed and $k$ being large. Observe that the EMC states that there should be a huge difference between the cases $s=2$ and $s\geq 3$: on the one hand $\frac{f(2k,k,2)}{\binom{2k}{k}}=\frac{1}{2}$ and $\frac{f(2k+1,k,2)}{\binom{2k+1}{k}}\rightarrow\frac{1}{2}$ as $k$ tends to infinity. On the other hand, $\frac{f(sk,k,s)}{\binom{sk}{k}}=\frac{s-1}{s}$, while $\frac{|{\mathcal{A}}_{k,s}|}{\binom{sk+1}{k}}\rightarrow(\frac{s-1}{s})^{2}$ and $\frac{|{\mathcal{B}}_{sk+1,k,s}|}{\binom{sk+1}{k}}\rightarrow 1-(1-\frac{1}{s})^{s-1}$ as $k$ tends to infinity, thus the EMC would yield an immediate drop in the limiting constant. Our main result shows that there is indeed a gap between $\frac{f(sk,k,s)}{\binom{sk}{k}}=\frac{s-1}{s}$ and $\frac{f(sk+1,k,s)}{\binom{sk+1}{k}}$.
Theorem 3.
For any $s\geq 3$ there exists a positive real $\varepsilon_{s}$ such that $f(sk+1,k,s)\leq\bigl{(}\frac{s-1}{s}-\varepsilon_{s}\bigr{)}\binom{sk+1}{k}$ holds for all $k$.
To prove Theorem 3, as an intermediate step, we will show the following stability version of Kleitman’s bound on $f(sk,k,s)$ with respect to the minimum degree.
Theorem 4.
There exist absolute constants $C$ and $\delta_{0}$ such that, if $s\geq 3$ and $\delta\leq\delta_{0}$, then any family ${\mathcal{F}}\subseteq\binom{[sk]}{k}$ with $\nu({\mathcal{F}})\leq s-1$ and minimum degree at least $\delta\binom{sk-1}{k-1}$ satisfies $|{\mathcal{F}}|\leq\bigl{(}\frac{s-1}{s}-\frac{(s-2)\delta}{s^{3}(s-1)C}\bigr{)}\binom{sk}{k}$.
2. Tools of the proofs: removal lemmas and shifting
In this section, we introduce some terminology and results from the literature that we will use in our proofs of Theorem 3 and Theorem 4. Removal lemmas are widely used in combinatorics, they state that if some combinatorial structure ${\mathcal{S}}$ contains only a few copies of some pattern $P$, then ${\mathcal{S}}$ can be made $P$-free while almost keeping the structure ${\mathcal{S}}$. We will need a removal lemma for intersecting families. A general result is the following.
Theorem 5 (Friedgut, Regev [16]).
Let $\gamma>0$ and $n,k$ positive integers with $\gamma n<k<(1/2-\gamma)n$. Then for any $\varepsilon>0$ there exists $\delta>0$ such that any ${\mathcal{F}}\subseteq\binom{[n]}{k}$ with at most $\delta|{\mathcal{F}}|\binom{n-k}{k}$ disjoint pairs can be made intersecting by removing at most $\varepsilon\binom{n-1}{k-1}$ sets from ${\mathcal{F}}$.
As we will point out in the last section, Theorem 5 could also be used to derive our results, but the following statement is better suited for our purposes.
Theorem 6 (Das, Tran [3]).
There exists an absolute constant $C$ such that if $n,k,\ell$ satisfy $n>2k\ell^{2}$, then for any ${\mathcal{F}}\subseteq\binom{[n]}{k}$ of size $(\ell-\alpha)\binom{n-1}{k-1}$ with at most $\bigl{(}\binom{\ell}{2}+\beta\bigr{)}\binom{n-1}{k-1}\binom{n-k-1}{k-1}$ disjoint pairs where $\max\{2\ell|\alpha|,|\beta|\}\leq\frac{n-2k}{(20C)^{2}n}$ there exists a family ${\mathcal{S}}$ that is a union of $\ell$ stars such that $|{\mathcal{F}}\triangle{\mathcal{S}}|\leq C\bigl{(}(2\ell-1)\alpha+2\beta\bigr{)}\frac{n}{n-2k}\binom{n-1}{k-1}$.
We state the case where $\ell=1,\alpha=0,n=sk>2k$ as a corollary, since this is what we will use in our proofs.
Corollary 7.
There exists an absolute constant $C$ such that for any ${\mathcal{F}}\subseteq\binom{[sk]}{k}$ of size $\binom{sk-1}{k-1}$ with at most $\frac{\beta}{s-1}\binom{sk-1}{k-1}\binom{(s-1)k}{k}$ disjoint pairs where $\beta\leq\frac{s-2}{s(20C)^{2}}$ there exists a star ${\mathcal{S}}$ such that $|{\mathcal{F}}\triangle{\mathcal{S}}|\leq\frac{2s}{s-2}C\beta\binom{sk-1}{k-1}$.
Consequently, if the maximum degree $\Delta({\mathcal{F}})$ of a family ${\mathcal{F}}\subseteq\binom{[sk]}{k}$ of size $\binom{sk-1}{k-1}$ is at most $(1-\delta)\binom{sk-1}{k-1}$ where $\delta\leq\frac{1}{200C}$, then ${\mathcal{F}}$ contains at least $\frac{\delta(s-2)}{2Cs(s-1)}\binom{sk-1}{k-1}\binom{(s-1)k}{k}$ disjoint pairs.
Let us now turn to our other tool in proving our results. The shifting operation was introduced by Erdős, Ko, and Rado [5] and is a very frequently used powerful tool in extremal set theory. For a set $F$ and family ${\mathcal{F}}$ of sets we define
$S_{i,j}(F)=\begin{cases}F\setminus\{j\}\cup\{i\},&\mbox{ if $i\notin F,j\in F,F\setminus\{j\}\cup\{i\}\notin F$;}\\ F,&\mbox{otherwise.}\end{cases}$
$S_{i,j}({\mathcal{F}})=\{S_{i,j}(F):F\in{\mathcal{F}}\}$. A family ${\mathcal{F}}\subseteq\binom{[n]}{k}$ is left-compressed if $S_{i,j}({\mathcal{F}})={\mathcal{F}}$ for all $1\leq i<j\leq n$. We will use the following lemma [6, Lemma 4.2(iv)].
Lemma 8 (Frankl [6]).
If ${\mathcal{F}}^{\prime}\subseteq\binom{[n]}{k}$ has matching number $\nu({\mathcal{F}})\leq s$, then there is a left-compressed ${\mathcal{F}}\subseteq\binom{[n]}{k}$ with $\nu({\mathcal{F}})\leq s$ and $\left|{\mathcal{F}}\right|=\left|{\mathcal{F}}^{\prime}\right|$.
We will need some more properties of left-compressed families. To state them, we need some definitions: for a family ${\mathcal{F}}\subseteq\binom{[n]}{k}$ and $x,y\in[n]$, we write ${\mathcal{F}}_{x}=\{F\in{\mathcal{F}}:x\in F\}$, ${\mathcal{F}}_{\overline{x}}=\{F\in{\mathcal{F}}:x\notin F\}$, ${\mathcal{F}}_{x,y}=\{F\in{\mathcal{F}}:x,y\in F\}$ and ${\mathcal{F}}_{x,\overline{y}}=\{F\in{\mathcal{F}}:x\in F,y\notin F\}$.
Lemma 9.
For any left-compressed family ${\mathcal{F}}\subseteq\binom{[n]}{k}$, we have
  1. (a)
    $\frac{n-k}{k}|{\mathcal{F}}_{n}|\leq|{\mathcal{F}}_{\overline{n}}|$,
  2. (b)
    $\frac{|{\mathcal{F}}_{n}|}{\binom{n-1}{k-1}}\leq\frac{|{\mathcal{F}}_{n-1,\overline{n}}|}{\binom{n-2}{k-1}}$.
Proof.
Inequality (a) follows from counting the pairs $(F,j)$ where $F\in{\mathcal{F}}_{n}$ and $j\not\in F$. The number of such pairs is $(n-k)|{\mathcal{F}}_{n}|$. Moreover, the set $S_{j,n}(F)$ is in ${\mathcal{F}}$ and any set $F\in{\mathcal{F}}_{\overline{n}}$ can be the $S_{j,n}$-image of at most $k$ sets in ${\mathcal{F}}_{n}$. Hence the number of such pairs is at most $k|{\mathcal{F}}_{\overline{n}}|$
To see (b), a similar double counting argument establishes $\frac{n-k}{k-1}|{\mathcal{F}}_{n,n-1}|\leq|{\mathcal{F}}_{n,\overline{n-1}}|$ and thus $|{\mathcal{F}}_{n}|=|{\mathcal{F}}_{n,n-1}|+|{\mathcal{F}}_{n,\overline{n-1}}|\leq\frac{n-1}{n-k}|{\mathcal{F}}_{n,\overline{n-1}}|$. Also, applying $S_{n-1,n}$, we obtain $|{\mathcal{F}}_{n,\overline{n-1}}|\leq|{\mathcal{F}}_{n-1,\overline{n}}|$. Therefore
$\frac{|{\mathcal{F}}_{n}|}{\binom{n-1}{k-1}}\leq\frac{\frac{n-1}{n-k}|{\mathcal{F}}_{n,\overline{n-1}}|}{\frac{n-1}{n-k}\binom{n-2}{k-1}}\leq\frac{|{\mathcal{F}}_{n-1,\overline{n}}|}{\binom{n-2}{k-1}}$
as claimed. ∎
3. Proofs
Proof of Theorem 4.
Let ${\mathcal{F}}\subseteq\binom{[sk]}{k}$ with $\nu({\mathcal{F}})\leq s-1$ and $\delta({\mathcal{F}})\geq\delta\binom{sk-1}{k-1}$. Let us write ${\mathcal{G}}=\binom{[sk]}{k}\setminus{\mathcal{F}}$ and consider an arbitrary subfamily ${\mathcal{G}}^{\prime}\subseteq{\mathcal{G}}$ with $|{\mathcal{G}}^{\prime}|=\frac{1}{s}\binom{sk}{k}=\binom{sk-1}{k-1}$. (By Theorem 1, we know $|{\mathcal{G}}|\geq\frac{1}{s}\binom{sk}{k}$, and so such a ${\mathcal{G}}^{\prime}$ exists.)
The minimum degree condition on ${\mathcal{F}}$ implies $\Delta({\mathcal{G}}^{\prime})\leq(1-\delta)\binom{sk-1}{k-1}$. Corollary 7 yields that if $\delta$ is sufficiently small, then ${\mathcal{G}}^{\prime}$ (and thus ${\mathcal{G}}$) contains at least $\frac{\delta(s-2)}{2Cs(s-1)}\binom{sk-1}{k-1}\binom{(s-1)k}{k}$ disjoint pairs.
Every set $G$ is contained in $M=\frac{1}{(s-1)!}\prod_{j=1}^{s-1}\binom{jk}{k}$ $s$-partitions of $[sk]$. Also, every disjoint pair $G,G^{\prime}$ is contained in $M^{\prime}=\frac{1}{(s-2)!}\prod_{j=1}^{s-2}\binom{jk}{k}$ $s$-partitions, and clearly one fixed $s$-partition contains $\binom{s}{2}$ disjoint pairs. Let us count the pairs $(G,\pi)$ with $\pi$ being an $s$-partition of $[sk]$ into $k$-sets, and $G\in{\mathcal{G}}$ being one of those $k$-sets. On the one hand the number of such pairs is $|{\mathcal{G}}|\cdot M$. On the other hand, by definition of ${\mathcal{F}}$ and ${\mathcal{G}}$, every $\pi$ contains at least one set of ${\mathcal{G}}$. Also, by the above observations, the number of $\pi$’s containing at least two sets from ${\mathcal{G}}$ is at least $\frac{\delta(s-2)}{2Cs(s-1)}\binom{sk-1}{k-1}\binom{(s-1)k}{k}\cdot\frac{M^{\prime}}{\binom{s}{2}}$. Putting these together,
$|{\mathcal{G}}|\cdot M\geq\frac{1}{s!}\prod_{j=1}^{s}\binom{jk}{k}+\frac{\delta(s-2)}{2Cs(s-1)}\binom{sk-1}{k-1}\binom{(s-1)k}{k}\cdot\frac{M^{\prime}}{\binom{s}{2}}.$
Dividing by $M$, we get $|{\mathcal{G}}|\geq\binom{sk-1}{k-1}+\frac{(s-2)\delta}{s^{2}(s-1)C}\binom{sk-1}{k-1}$ and thus
$|{\mathcal{F}}|=\binom{sk}{k}-|{\mathcal{G}}|\leq\binom{sk-1}{k}-\frac{(s-2)\delta}{s^{2}(s-1)C}\binom{sk-1}{k-1}=\Bigl{(}\frac{s-1}{s}-\frac{(s-2)\delta}{s^{3}(s-1)C}\Bigr{)}\binom{sk}{k}$
as claimed. ∎
Proof of Theorem 3.
Let ${\mathcal{F}}\subseteq\binom{[sk+1]}{k}$ be a family of sets with $\nu({\mathcal{F}})<s$, $|{\mathcal{F}}|=f(sk+1,k,s)$ and let us write $n=sk+1$.
By Lemma 8, we can assume that ${\mathcal{F}}$ is left-compressed. By Theorem 1, we have $|{\mathcal{F}}_{\overline{n}}|\leq\frac{s-1}{s}\binom{sk}{k}$. Let $\delta_{0}$, $C$ be as in Theorem 4, and let $\varepsilon^{*}_{s}=\min\bigl{\{}\frac{(s-2)\delta_{0}}{s^{3}(s-1)C},\frac{s-1}{s}-\delta_{0}\bigr{\}}$.
Suppose first $|{\mathcal{F}}_{n}|\leq\bigl{(}\frac{s-1}{s}-\varepsilon^{*}_{s}\bigr{)}\binom{sk}{k-1}$, then
$|{\mathcal{F}}|=|{\mathcal{F}}_{\overline{n}}|+|{\mathcal{F}}_{n}|\leq\frac{s-1}{s}\binom{sk}{k}+\Bigl{(}\frac{s-1}{s}-\varepsilon^{*}_{s}\Bigr{)}\binom{sk}{k-1}\leq\Bigl{(}\frac{s-1}{s}-\frac{\varepsilon^{*}_{s}}{s+1}\Bigr{)}\binom{sk+1}{k}.$
On the other hand, if $|{\mathcal{F}}_{n}|\geq\bigl{(}\frac{s-1}{s}-\varepsilon^{*}_{s}\bigr{)}\binom{sk}{k-1}$, then we use the fact that left-compression implies $\delta({\mathcal{F}}_{\overline{n}})=|{\mathcal{F}}_{n-1,\overline{n}}|$ and Lemma 9(b) to conclude that
$\frac{\delta({\mathcal{F}}_{\overline{n}})}{\binom{sk-1}{k-1}}=\frac{|{\mathcal{F}}_{n-1,\overline{n}}|}{\binom{sk-1}{k-1}}\geq\frac{|{\mathcal{F}}_{n}|}{\binom{sk}{k-1}}\geq\frac{s-1}{s}-\varepsilon^{*}_{s}.$
So one can apply Theorem 4 with $\delta=\delta_{0}$ and since $\frac{s-1}{s}-\epsilon_{s}^{*}\geq\delta_{0}$, it gives $|{\mathcal{F}}_{\overline{n}}|\leq\bigl{(}\frac{s-1}{s}-\varepsilon^{*}_{s}\bigr{)}\binom{sk}{k}$. According to Lemma 9(a), we have $|{\mathcal{F}}_{n}|\leq\frac{k}{(s-1)k+1}|{\mathcal{F}}_{\overline{n}}|$ and thus
$|{\mathcal{F}}|=|{\mathcal{F}}_{n}|+|{\mathcal{F}}_{\overline{n}}|\leq\Bigl{(}\frac{s-1}{s}-\varepsilon^{*}_{s}\Bigr{)}\binom{sk+1}{k}.$
So $\varepsilon_{s}=\varepsilon^{*}_{s}/(s+1)$ works in both cases. ∎
4. Concluding remarks
As we mentioned in Section 2, Theorem 3 and Theorem 4 can be proved using the “$\varepsilon$-$\delta$ removal lemma” of Theorem 5 with the help of Frankl’s result [7] on the maximum size of intersecting families with bounded maximum degree. Also, there is a large literature on the size of intersecting families with constraints on some other parameters: minimum degree [21] and diversity [11, 13, 24, 25]. Supersaturation results on the minimum number of disjoint pairs in a family are also known [2]. Although, exact results are usually known for values of $n$ large enough compared to $k$, weaker bounds could be turned into bounds on $f(n,k,s)$ with the methods used in our note.
The authors would like to acknowledge the support of Iowa State University, where this work was done while the second author was a visitor.